OpenStudy (anonymous):

A convex heptagon has interior angles that measure xo, 160o, 150o, 2xo, 150o, 160o, and 2xo. What is the value of x?

OpenStudy (anonymous):

would u pls draw ?

OpenStudy (tfht):

so to find the total degrees in a figure you do (the # of sides-2)*180 so it would be (7-2)*180

OpenStudy (tfht):

so if you simplify it, you get 900degrees

OpenStudy (tfht):

so then you subtract 160,150, 150, and 160 to get 280. then you add x+2x which equals 3x then you divide 280/3 to get that x=93 (1/3)

OpenStudy (tfht):

and there you go

OpenStudy (tfht):



OpenStudy (tfht):

and i realized that i messed up so it would be 5x=280 then you divide both sides by 5 to get x=56 sorry my bad!!!

OpenStudy (tkhunny):

I never remember the sum of interior angles. It's trivial to remember the sum of exterior angles. It's 360º EVERY TIME. If it's a regular heptagon, that leaves 180º - 360º/7 = 128º + (4/7)º There are 7 of them, so 900º for the interior angles. I do recommend remember the faster way to the sum or interior angles, but if it slips your mind, there is always a backup plan.

OpenStudy (tfht):

or you could use the formula that i gave you...